Module: HOB+0107/16000 (Backup: BKG+0142/4000)
Data volume at beginning: 8318
- Fs complained 've' was not in remote at start of experiment. Switching to remote resolved.
- Attenuation recommended by notes was too low. Changed to 6db in IFD as per CRDS73 prc file.
- 02:15 UT: He supply pressure has shot up from <285 psi to 310 psi because Brett Reid introduced new supply (Ross).